# Social, Legal, Ethical, Empathetic, and Cultural Rules: Compilation and Reasoning (Extended Version)

### v1

@@ 1 Introduction
The rise of AI-based and autonomous systems is posing concerns and causing apprehension due to potential negative repercussions arising from their behavior or decisions. Philosophers, engineers, and legal experts are actively exploring avenues to regulate the actions of these systems, which are becoming increasingly integrated into our daily existence. One way to exert control over these systems involves the identification of rules dictating their conduct during collaboration and interaction with humans.
Normative multi-agent systems (Chopra et al. 2018) are systems where the behaviour of the participants is affected by the norms established therein. Townsend et al. (2022) propose a methodology to elicit normative rules for multi-agent systems. The process follows an iterative approach which is:
@@
In the process, philosophers, ethicists, lawyers, domain experts, and other professionals can extract normative rules from a specific domain. These rules are referred to as SLEEC rules, which stands for social, legal, ethical, empathetic, and cultural rules. These rules are aimed to play a crucial role for facilitating the creation of tools and instruments to engineer systems that conform adequately to the established norms, while also aiding in the verification of existing system behavior. However, the SLEEC rules obtained through elicitation are typically expressed in natural language. Therefore, to be employed effectively in AI systems, these rules need to be translated into a machine-interpretable language. Fortunately, the elicitation process generates rules that are characterized by a clear pattern.
@@ 2.1 Propositional Logic
A positive literal is a propositional variable $p$ and a negative literal is the negation of a propositional variable $\lnot p$. A clause is a disjunction of literals. A formula in CNF is a conjunction of clauses. A formula in 3CNF is a conjunction of clauses of three literals. A Horn formula is a conjunction of clauses containing at most one negative literal.
@@ 3.2 The Logical Form of ‘Unless’
$v(\phi\operatorname{{\sf UNLESS}}\psi)=true\text{ iff }v(\lnot\phi)=false\text{ or }v(\psi)=true.$
@@ 7.1 Propositional Reasoning
The SLEEC rule of Example 1 is compiled into the following propositional formula: [^4] Any off-the-shelf SAT solver can be used: http://www.satlive.org/solvers/. . $(a\land\lnot\lnot d\rightarrow o)\land(a\land\lnot d\land\lnot h\rightarrow n\land s)\land(a\land\lnot d\land h\rightarrow o)$. Let us note it $\phi_{SLEEC}$.
@@
To ensure the soundness of our solution, it is imperative that it does not propose any spurious obligations. Fortunately, our solution meets this requirement, as no additional obligations are entailed by the previous formulas. For instance, $(a\land d\land h)\land\phi_{SLEEC}\not\models n\lor s$, because the interpretation $v=\{a\mapsto true,d\mapsto true,h\mapsto true,\mathbf{n\mapsto false},o\mapsto true,\mathbf{s\mapsto false}\}$ is a model of $(a\land d\land h)\land\phi_{SLEEC}$. Similarly, we have that $(a\land\lnot d\land\lnot h)\land\phi_{SLEEC}\not\models o$, because the interpretation $v^{\prime}=\{a\mapsto true,d\mapsto false,h\mapsto false,n\mapsto true,\mathbf{o\mapsto false},s\mapsto true\}$ is a model of $(a\land\lnot d\land\lnot h)\land\phi_{SLEEC}$.
@@ 7.2 Logic Programming: Answer Set Programming
We have chosen to represent a situation where the user is not dressed (d(user) is not present) but highly distressed (h(user)). The clause d(someoneelse) is present for technical reason, because all predicates must appear in the head of a rule. We obtain the expected model in clingo (5.7.0):555 https://potassco.org/clingo/run/ can be used for testing.
@@ 7.3 Logic Programming: PROLOG
a(user,curtains).  % user asks the robot to open the curtains
@@
To get all the obligations, we can ask the PROLOG query: [^6] https://swish.swi-prolog.org/ can be used for testing.
@@ 8 Conclusions
The ethical considerations are made during the elicitation phase to obtain SLEEC rules that are adequate for an application (Townsend et al. 2022). SLEEC rules prevent the robot’s behaviours that could do harm, violating the moral principles. Our contribution allows the unambiguous and rigorous implementation of SLEEC rules that participate in the specification of the actual robot behaviour. However, an ethical review of the system as a whole remains necessary.
@@ Acknowledgements
The work was partially supported by the PRIN project P2022RSW5W RoboChor: Robot Choreography, the PRIN project 2022JKA4SL HALO: etHical-aware AdjustabLe autOnomous systems, and of the MUR (Italy) Department of Excellence 2023–2027.
@@ Appendix A Proof of Proposition 2
Every conjunction in every SLEEC rule compile according to Definition 4 is of the form $C_{0}\land\ldots C_{i}\land\lnot C_{i+1}\rightarrow O_{i}$ or $C_{0}\land\ldots C_{i}\land C_{n}\rightarrow O_{n}$.
The first form is equivalent to $\lnot C_{0}\lor\ldots\lnot C_{i}\lor C_{i+1}\lor O_{i}$. Since $C_{0},\ldots C_{i+1}$ are negative literals and $O_{i}$ is a positive literal, it is a Horn clause, with $C_{i+1}$ being the only negative literal ($\lnot C_{0},\ldots,\lnot C_{i}$ are positive literals).
The second form is also equivalent to a Horn clause (with only positive literals).
$\Gamma$ entails $o$ iff the set of Horn clauses composed of all the clauses in $\Gamma$ and the Horn clause $\lnot o$ is not satisfiable. Using 5, we know that this can be decided in polynomial time, because HORNSAT is in PTIME (Dowling and Gallier 1984).
@@ Appendix B Another Scenario: Post Office
This scenario makes use of more intricate relations between the objects of the domain. In fact, this scenario is more aptly represented with a predicative language like PROLOG, than with plain propositional logic.
An administrative service requires to come to a public desk, and requires to check-in at the time of arrival. At check-in the user is nominatively assigned a number, that is incremented. The default SLEEC rule for the on-site scheduling is simply as follows.
@@
A user can specify soft-ethics rules. E.g., a user can prefer to swap her turn with another user who checked in later, if this user has some kind of handicap; and a user with a certain handicap can prefer to swap her turn with another user who checked earlier. The on-site scheduling automaton is capable of recognising the handicaps of the users, and is capable of accessing their soft ethics. It must then take all this information into account when calling the next user. The SLEEC rule becomes as follows.

### v3

@@ 1 Introduction
The rise of AI-based and autonomous systems is posing concerns and causing apprehension due to potential negative repercussions arising from their behavior or decisions. Philosophers, engineers, and legal experts are actively exploring avenues to regulate the actions of these systems, which are becoming increasingly integrated into our daily existence. One way to exert control over these systems involves the identification of rules that dictate their conduct during collaboration and interaction with humans. [^1] This paper was originally published as Troquard et al. (2024).
Normative multi-agent systems (Chopra et al. 2018) are systems where the behaviour of the participants is affected by the norms established therein. Townsend et al. (2022) propose a methodology to elicit normative rules for multi-agent systems. The process follows an iterative approach:
@@
In the process, philosophers, ethicists, lawyers, domain experts, and other professionals can extract normative rules from a specific domain. These rules are referred to as SLEEC rules, which stands for social, legal, ethical, empathetic, and cultural rules. These rules are aimed to play a crucial role in facilitating the creation of tools and instruments to engineer systems that conform adequately to the established norms, while also aiding in the verification of existing system behavior. However, the SLEEC rules obtained by elicitation are typically expressed in natural language. Therefore, to be employed effectively in AI systems, these rules need to be translated into a machine-interpretable language. Fortunately, the elicitation process generates rules that are characterized by a clear pattern.
@@ 2.1 Propositional Logic
A positive literal is a propositional variable $p$ and a negative literal is the negation of a propositional variable $\lnot p$. A clause is a disjunction of literals. A formula in CNF is a conjunction of clauses. A formula in 3CNF is a conjunction of clauses of three literals.
@@ 3.2 The Logical Form of ‘Unless’
$v(\phi\operatorname{{\sf UNLESS}}\psi)=true\text{ iff }v(\lnot\psi)=false\text{ or }v(\phi)=true.$
@@ 7.1 Propositional Reasoning
The SLEEC rule of Example 1 is compiled into the following propositional formula: $(a\land\lnot\lnot d\rightarrow o)\land(a\land\lnot d\land\lnot h\rightarrow n\land s)\land(a\land\lnot d\land h\rightarrow o)$. Let us note it $\phi_{SLEEC}$. Any off-the-shelf SAT solver can now be used for reasoning.
@@
To ensure the soundness of our solution, it is imperative that it does not propose any spurious obligations. Fortunately, our solution meets this requirement, as no additional obligations are entailed by the previous formulas. For instance, $(a\land d\land h)\land\phi_{SLEEC}\not\models n\lor s$, because the interpretation $v=\{a\mapsto true,d\mapsto true,h\mapsto true,\bm{n\mapsto false},o\mapsto true,\bm{s\mapsto false}\}$ is a model of $(a\land d\land h)\land\phi_{SLEEC}$. Similarly, we have that $(a\land\lnot d\land\lnot h)\land\phi_{SLEEC}\not\models o$, because the interpretation $v^{\prime}=\{a\mapsto true,d\mapsto false,h\mapsto false,n\mapsto true,\bm{o\mapsto false},s\mapsto true\}$ is a model of $(a\land\lnot d\land\lnot h)\land\phi_{SLEEC}$.
@@ 7.2 Logic Programming: Answer Set Programming
We have chosen to represent a situation where the user is not dressed (d(user) is not present) but highly distressed (h(user)). The clause d(someoneelse) is present for technical reason, because all predicates must appear in the head of a rule. We obtain the expected model in the ASP system clingo (5.7.0):
@@ 7.3 Logic Programming: PROLOG
a(user,curtains).  % user asks robot open the curtains
@@
To get all the obligations, we can ask the PROLOG query:
@@ 8 Conclusions
The ethical considerations are addressed during the elicitation phase to derive SLEEC rules that are suitable for an application (Townsend et al. 2022). SLEEC rules are designed to prevent the robot’s behaviors that could potentially cause harm, thereby avoiding violations of moral principles. Our contribution enables the unambiguous and rigorous implementation of SLEEC rules, contributing to the specification of the actual robot behavior. Nevertheless, a comprehensive ethical review of the entire system remains essential.
@@ Acknowledgements
The work was partially supported by the PRIN project P2022RSW5W RoboChor: Robot Choreography, the PRIN project 2022JKA4SL HALO: etHical-aware AdjustabLe autOnomous systems, and by the MUR (Italy) Department of Excellence 2023–2027.
@@ Appendix A Proof of Proposition 2
A Horn formula is a conjunction of clauses containing at most one positive literal. A dual-Horn formula is a conjunction of clauses containing at most one negative literal.
@@
Every conjunction in every SLEEC rule compiled according to Definition 4 is of the form $C_{0}\land\ldots C_{i}\land\lnot C_{i+1}\rightarrow O_{i}$ or $C_{0}\land\ldots C_{i}\land C_{n}\rightarrow O_{n}$.
The first form is equivalent to $\lnot C_{0}\lor\ldots\lnot C_{i}\lor C_{i+1}\lor O_{i}$. Since $C_{0},\ldots C_{i+1}$ are negative literals and $O_{i}$ is a positive literal, it is a dual-Horn clause, with $C_{i+1}$ being the only negative literal ($\lnot C_{0},\ldots,\lnot C_{i}$ are positive literals).
The second form is also equivalent to a dual-Horn clause (with only positive literals).
$\Gamma$ entails $o$ iff the set of dual-Horn clauses composed of all the clauses in $\Gamma$ and the dual-Horn clause $\lnot o$ is not satisfiable. We obtain an equi-satisfiable set of Horn clauses by simply negating all variables. Using 5, we know that this reduction can be done in polynomial time. Hence, we know that the problem can be decided in polynomial time because HORNSAT is in PTIME (Dowling and Gallier 1984).
@@ Appendix B Another Scenario: Post Office
This scenario makes use of more intricate relations among the objects of the domain. In fact, this scenario is more aptly represented with a predicative language like PROLOG, than with plain propositional logic.
An administrative service requires to come to a public desk, and requires to check-in at the time of arrival. There is an integer counter. At check-in, the user is assigned a number—the value of the counter—and the counter is then incremented. The default SLEEC rule for the on-site scheduling is simply as follows.
@@
A user can specify preferences or soft ethics (Floridi 2018). E.g., a user can prefer to swap her turn with another user who checked in later, if this user has some kind of handicap; and a user with a certain handicap can prefer to swap her turn with another user who checked in earlier. The on-site scheduling automaton is capable of recognising the handicaps of the users, and is capable of accessing their soft ethics. It must then take all this information into account when calling the next user. The SLEEC rule becomes as follows.